How much interest will be added to Php75,000 at 2.25% due at the end of 9 months?

1
Expert's answer
2021-11-21T16:51:51-0500

"I=\frac {PRT} {100}"


"I=\frac {75,000×2.25×0.75}{100}"


"I=\frac {126,562.5} {100}"


"I=Php1265.63"